What potential difference is needed to give a helium nucleus (Q=3.2×10^−19 C) 120 keV of kinetic energy?
Answer #1

Kinetic energy = potential energy

120keV =QV

V = \frac{120\times 10^{-3}(1.6\times 10^{-19}C)V}{3.2\times 10^{-19}C}=60\times 10^{-3}V

Potential difference = 0.06V